# UniRestore: Unified Perceptual and Task-Oriented Image Restoration Model Using Diffusion Prior

### v1

@@ 2.1 Perceptual Image Restoration
Perceptual image restoration aims to enhance the visual quality of images as perceived by humans, and it can be categorized into single degradation and multiple degradation restoration. Early work in single degradation restoration, such as SRCNN [13], focused on specific degradations to improve image quality, leading to significant advancements in super-resolution [26, 33, 79], denoising [74, 8], dehazing [6, 74], deraining [59, 69], low-light enhancement [72, 78], and deblurring [41, 47]. To tackle multiple degradations simultaneously, methods like MPRNet [74] and NAFNet [2] introduced unified solutions. Recently, transformer-based approaches, such as SwinIR [31] and Restormer [73], have gained traction for their versatility. Additionally, holistic approaches like All-in-One [30], TransWeather [58], and PromptIR [45] have focused on improving visual quality across a wide range of conditions while providing enhanced adaptability and performance. While these methods excel in enhancing visual quality, they do not always guarantee improved performance in downstream vision tasks.
@@ 2.2 Task-oriented Image Restoration
Task-oriented image restoration aims to enhance downstream task performance, as studies [27, 4, 71] show that image degradation significantly impairs high-level task accuracy. DDP [64] aligns feature representations between low- and high-quality images to improve classification accuracy. SFDUnet [70] employs self-feature distillation and uncertainty modeling to extract high-quality-like features from degraded images, enhancing recognition in challenging regions. URIE [55] integrates image enhancement and recognition tasks in an end-to-end framework to mitigate degradation effects. DIP [37] adapts image processing dynamically based on degradation factors for better recognition outcomes. VDR-IR [71] unifies semantic representations of diverse degraded images to recover intrinsic semantics effectively. While these methods enhance downstream task performance, they may result in images that are less visually pleasing.
@@ 2.3 Diffusion Model
Diffusion Models (DMs) [19] leverage a parameterized Markov chain to optimize the lower variational bound on the likelihood function, achieving state-of-the-art results in sample quality[48, 56, 57, 24] and various applications [67, 53, 81]. ControlNet [77] further refines this architecture by optimizing it for conditional image generation tasks. It incorporates conditions such as segmentation maps or textual prompts directly into the diffusion process through a specialized encoder, which modifies the latent representation of the input, enabling ControlNet to generate images that closely align with specified conditions [80, 75].
Recently, DMs have become integral to image restoration, including tasks such as super-resolution [49, 29], inpainting [39, 48, 10], and degradation restoration [43, 83, 66]. StableSR [60] effectively leverages diffusion priors for real-world super-resolution, resulting in superior reconstruction quality. DiffUIR [82] utilizes a specialized hourglass architecture to map degraded inputs to high-quality outputs, enhancing both global and local features effectively. DiffBIR [36] follows a two-stage restoration strategy, first addressing specific degradations and then refining the image quality through a diffusion generation model. Despite these advances, adopting pre-trained diffusion models for both PIR and TIR remains an open challenge.
@@ 3 Proposed Method
Figure 2: Overview of UniRestore. UniRestore augments the diffusion model by integrating a CFRM and a TFA. The training process is divided into two stages: In the first stage, CFRM, Controller, and SC-Tuner are trained to restore clear encoder and latent features. In the second stage, the TFA is trained to adapt the restored encoder features and latent features for various downstream tasks, using task-specific prompts at the decoder to control the output restoration.
@@ 3.2 Complementary Feature Restoration Module
The objective of the CFRM is to restore and enhance features within the encoder, thereby providing complementary inputs to the decoder. As shown in  Figure 3, the CFRM is integrated into the output of each encoder layer and consists of four steps:
Feature Enhancement:The enhancement begins with a NAFBlock [9], followed by a convolutional layer and group normalization [65]. Input features have dimensions $(C^{{}^{\prime}},H,W)$, where $C^{{}^{\prime}}$ denotes the number of channels, $H$ the height, and $W$ the width. These features are expanded fourfold to dimensions $(4C^{{}^{\prime}},H,W)$ and subsequently divided into $l$ groups, resulting in $(C^{{}^{\prime\prime}},H,W)$, where $C^{{}^{\prime\prime}}=4C^{{}^{\prime}}/l$.
@@ 4 Implementation Details
4.1 Training Dataset
4.2 Evaluation Dataset
4.3 Evaluation Protocol
For the evaluation of PIR, we utilize multiple metrics, including Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM). For image classification, we measure performance using accuracy (ACC), and for semantic segmentation, we use the mean Intersection over Union (mIoU).
@@ 5.3 Task-oriented Image Restoration
Semantic Segmentation. In the training stage, we adopt DeepLabv3+ [3] as the segmentation model, while for evaluation, we employ both DeepLabv3+ [3] and RefineNet-lw [42]. Both models are pretrained on the training set of Cityscapes dataset [11]. As shown in Table 3, UniRestore achieves decent performance in semantic segmentation on both seen datasets and the unseen dataset. These results underscore UniRestore’s ability to restore fine-grained details crucial for semantic segmentation tasks, demonstrating that the TFA modules integrate diffusion features with restored features to produce restored images with high-quality feature representations. Similar to the results observed in classification, PIR methods show limited performance in semantic segmentation compared to TIR methods, with performance decreasing in multiple downstream task scenarios. A qualitative comparison in Figure 6 shows that images restored by UniRestore enable segmentation models to generate more accurate object boundaries in the segmented results.

### v2

@@ Abstract
Project page: https://unirestore.github.io/* Indicates equal contribution.
@@ 2.1 Perceptual Image Restoration
Perceptual image restoration aims to enhance the visual quality of images as perceived by humans, and it can be categorized into single degradation and multiple degradation restoration. Early work addressed single degradation such as denoising [72, 8], dehazing [13, 6, 72], deraining [57, 67], low-light enhancement [70, 75], and deblurring [38, 44]. To tackle multiple degradations, methods like MPRNet [72] and NAFNet [2] introduced unified solutions, while transformer-based approaches such as SwinIR [29] and Restormer [71] gained traction for their versatility. Additionally, holistic approaches like All-in-One [28], TransWeather [55] and PromptIR [42] focused on improving visual quality across a wide range of conditions while providing enhanced adaptability and performance. While these methods excel in enhancing visual quality, they do not always improve downstream vision tasks.
@@ 2.2 Task-oriented Image Restoration
Task-oriented image restoration aims to enhance downstream task performance, as studies [25, 4, 69, 62, 68] show that image degradation significantly impairs high-level task accuracy. URIE [52] integrates image enhancement and recognition in an end-to-end framework to mitigate degradation effects. DIP [34] adapts image processing dynamically based on degradation factors. VDR-IR [69] unifies semantic representations of diverse degraded images to recover intrinsic semantics effectively. While these methods enhance downstream task performance, they may result in images that are less visually pleasing.
@@ 2.3 Diffusion Model
Diffusion Models (DMs) [18] leverage a parameterized Markov chain to optimize the lower variational bound on the likelihood function, achieving state-of-the-art results in sample quality [45, 53, 54, 23] and various applications [65, 50, 76, 74]. Recently, DMs have become integral to image restoration, including super-resolution [46, 27], inpainting [36, 45, 10], and degradation restoration [40, 78, 64]. StableSR [58] effectively leverages diffusion priors for real-world super-resolution, resulting in superior reconstruction quality. DiffUIR [77] utilizes a specialized hourglass architecture to map degraded inputs to high-quality outputs, enhancing both global and local features effectively. DiffBIR [33] follows a two-stage restoration strategy, first addressing specific degradations and then refining the image quality through a diffusion generation model. Despite these advances, adopting pre-trained diffusion models for both PIR and TIR remains an open challenge.
@@ 3 Proposed Method
Figure 2: Overview of UniRestore. UniRestore augments the diffusion model by incorporating CFRMs and TFAs within the pre-trained autoencoder. The training process is divided into two stages: In the first stage, CFRM, Controller, and SC-Tuner are trained to restore clear encoder and latent features. In the second stage, the TFA is trained to adapt the restored encoder features and latent features for various downstream tasks, using task-specific prompts at the decoder to control the output restoration.
@@ 3.2 Complementary Feature Restoration Module
The CFRM aims to restore and enhance features within the encoder, thereby providing complementary inputs to the decoder. As shown in  Figure 3, the CFRM is integrated into the output of each encoder layer and consists of four steps:
Feature Enhancement:The enhancement begins with a NAFBlock [9], followed by a convolutional layer and group normalization [63]. Input features have dimensions $(C^{{}^{\prime}},H,W)$, where $C^{{}^{\prime}}$ denotes the number of channels, $H$ the height, and $W$ the width. These features are expanded fourfold to dimensions $(4C^{{}^{\prime}},H,W)$ and then divided into $l$ groups, resulting in $(C^{{}^{\prime\prime}},H,W)$, where $C^{{}^{\prime\prime}}=4C^{{}^{\prime}}/l$.
@@ 4 Implementation Details
4.1 Training Dataset
4.2 Evaluation Dataset
4.3 Evaluation Protocol
For the evaluation of PIR, we utilize Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM). For image classification, we measure performance using accuracy (ACC), and for semantic segmentation, we use the mean Intersection over Union (mIoU).
@@ 5.3 Task-oriented Image Restoration
Semantic Segmentation. In the training stage, we adopt DeepLabv3+ [3] as the segmentation model, while for evaluation, we employ both DeepLabv3+ [3] and RefineNet-lw [39]. Both models are pretrained on the Cityscapes [11] training set. As shown in Table 3, UniRestore achieves decent performance in semantic segmentation on both seen datasets and the unseen dataset. These results underscore UniRestore’s ability to restore fine-grained details crucial for semantic segmentation tasks, demonstrating that the TFA modules integrate diffusion features with restored features to produce restored images with high-quality feature representations. Similar to the results observed in classification, PIR methods show limited performance in semantic segmentation compared to TIR methods, with performance decreasing in multiple downstream task scenarios. A qualitative comparison in Figure 6 shows that images restored by UniRestore enable segmentation models to generate more accurate object boundaries in the segmented results.
5.4 Ablation Study
@@ 6 Conclusion
Acknowledgements
This research was supported by Taiwan’s National Science and Technology Council under Grant NSTC 111-2221-E-002-136-MY3 and the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) contract number 140D0423C0074. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.